OpenStudy (anonymous):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)?

OpenStudy (anonymous):

Plug the vertex and one of the roots into the standard form, then solve for the missing variable.

OpenStudy (anonymous):

I have done it several times and I am not getting the right equation.

OpenStudy (anonymous):

Hm, what do you get when you substitute the vertex in?

OpenStudy (anonymous):

I tried it a few different ways and once I got the final equation as f(x)=4x^2-8x+12 and another time I got f(x)=2x^2+2x-6 neither are right when I check the graph.

OpenStudy (anonymous):

@ig0p0g0

OpenStudy (anonymous):

Try y=a(x-h)^2 +k, where h,k is the vertex. What do you get when you plug in the vertex?

OpenStudy (anonymous):

y=a(x-(-1))^2 + -8?

OpenStudy (anonymous):

Yep...probably the easiest next step is to plug in one of the roots. What do you get?

OpenStudy (anonymous):

y=a(-1 - (-1))^2 +8?

OpenStudy (anonymous):

What's the value of y if that's a root?

OpenStudy (anonymous):

the other root?

OpenStudy (anonymous):

Think about...what does it mean for something to be a root?

OpenStudy (anonymous):

where it hits the line on the x axis

OpenStudy (anonymous):

@ig0p0g0

OpenStudy (anonymous):

Yes! What's the y value there?

OpenStudy (anonymous):

-8? @ig0p0g0

OpenStudy (anonymous):

The y value of a point on the x axis....

OpenStudy (anonymous):

....any point on the x axis...

OpenStudy (anonymous):

ok any one at all

OpenStudy (anonymous):

(4,0) is on the x-axis. (4,3) is not. There's something you know about any point on the x-axis.

OpenStudy (anonymous):

no y valuee

OpenStudy (anonymous):

well, zero for y.

OpenStudy (anonymous):

So go back to this... y=a(-1 - (-1))^2 +8? and plug in zero for y.

OpenStudy (anonymous):

0=a(-1 - (-1))^2 +8

OpenStudy (anonymous):

OK, what's a?

OpenStudy (anonymous):

Ooo...something's wrong there...

OpenStudy (anonymous):

Oh, the root is 1, not -1

OpenStudy (anonymous):

So you want 0=a(1 - (-1))^2 +8

OpenStudy (anonymous):

Gotta go but you're almost done.